# Harvesting Private Medical Images in Federated Learning Systems with Crafted Models

### v1

@@ III Threat Model
In Fig.1 we demonstrate the threat model of our attack. We consider the parameter server $S$ to be a malicious party that is curious about the training samples held by clients (e.g., private patient information). We consider the parameter server to be a proactive attacker and can actively modify model parameters and architectures from $G$ to $\hat{G}$ to achieve the attack goals, following the same assumption as [25, 24, 37]. We assume the communication channels between the server and clients are secured and all messages can be authenticated. We assume the state-of-the-art SA protocols are in place and the server can only get access to the aggregated model updates $\sum_{i=1}^{n}\delta_{i}$ without knowing anything about the individual values $\delta_{i}$. We consider the attacker is able to collect or obtain an auxiliary dataset $D_{aux}$ that has the same data format and can represent the target dataset well. The attacker can leverage various online resources such as publicly available datasets, image searching tools, and image generative tools to fulfill this requirement. For our use case, the availability of public chest X-ray datasets makes this task trivial. The goal of the attacker is to recover the local data samples of a target client $c_{target}$ from the aggregated model updates $\sum_{i=1}^{n}\delta_{i}$. This can be mathematically expressed as: $D_{target}=Reverse(\sum_{i=1}^{n}\delta_{i},\hat{G})$.
IV Attack Method
IV-A Attack Overview
@@
We decompose the complex recovery problem into two different tasks including the individual model update retrieval and efficient model update reversion. To accomplish them, we require the attacker to place an additional two-linear-layer module $L_{adv}$ with the rectified linear unit (ReLU) activation function in between at the beginning of the original global model, i.e. $\hat{G}=G\oplus L_{adv}$. The dimension of this module is identical to the image dimension. We require the attacker to initialize the parameters of the linear module $L_{adv}$ to form different adversarial modules and distribute them to different clients accordingly. For all the other clients except the victim, the attacker crafts the “zero gradient” modules $L_{zero}$ to ensure that the gradients and model updates of these clients are always zero, i.e. $\hat{G}_{others}=G\oplus L_{zero}$. By doing this, the attacker guarantees that only the model update of the victim client is exposed, accomplishing task one. For the victim client, the attacker crafts a “linear leakage” module $L_{linear}$, aiming to reverse its model update back to local training samples efficiently, i.e. $\hat{G}_{target}=G\oplus L_{linear}$. This module requires an auxiliary dataset to help generate essential parameters and can ensure that samples are perfectly recovered with a mathematical proof, accomplishing task two. Detailed attack flow and module designs are introduced in the next section.
IV-B Detailed Attack Flow
@@
We demonstrate the attack flow in Fig.2. MediLeak is a two-phase attack including 1) attack preparation and 2) sample reconstruction phases. In the attack preparation phase, the attacker crafts the adversarial global model $\hat{G}$ and initializes it with different model parameters including $L_{zero}$ and $L_{linear}$ (step ①) before publishing them to different clients (step ②). Then in the second phase, the attacker collects the aggregated model updates and uses an analytical method to reverse it back to local training samples (step ③).
@@
In this phase, the attacker crafts both the “linear leakage” module and “zero gradient” module once at the outset. Both modules require the estimation of essential parameters of a representative auxiliary dataset $D_{aux}$. In detail, the attacker first estimates the cumulative density function (CDF) of the brightness feature $h(x)$ of the auxiliary dataset $D_{aux}$, denoted by $\psi(h(x))$, to represent the CDF of the local training dataset (which is unavailable), where $x$ refers to the input vector. After that, the attacker divides the distribution $\psi$ into equally $k$ bins by calculating $h_{j}=\psi^{-1}(j/k)$ where $j\in\{1,2,\cdots,k\}$, $\psi^{-1}$ refers to the inverse function of $\psi$, and $k$ equals to the neuron number of the first linear layer. By doing so, the brightness of a random input vector $x$ denoted by $h(x)$ will have the same probability of falling into each bin. This bin vector $\mathbf{H}=\left[h_{1},h_{2},\cdots,h_{k}\right]$ is the key vector to craft the model parameters of the adversarial global model $\hat{G}$ to form both attack modules.
@@
Zero Gradient Module: The “zero gradient” module is initialized in the same way as the “linear leakage” module except in step (2), in which the attacker has the bias vector $b_{1}$ equal to $\mathbf{H^{\prime}}=-\left[h_{k},h_{k},\cdots,h_{k}\right]$. By doing so, the output of the first linear layer will always be smaller than zero because $h_{k}$ is the largest possible brightness and all possible input $x$’s brightness is smaller than it. Considering we use the ReLU activation function after the first layer, the input to the second layer and the gradients of the first layer shall always be zero because of the ReLU function’s mathematical property. This results in zero model updates for all clients except the target victim. Therefore, the aggregated model updates are identical to the model updates of the victim client, i.e. $\sum_{i=1}^{n}\delta_{i}=\mathbf{0}+\mathbf{0}+\cdots+\mathbf{0}+\delta_{target}=\delta_{target}$, exposing the model updates of the target victim.
Figure 3: Recovered examples from the COVIDx CXR-4 dataset. The original images are shown in the first row and the recovered ones are shown in the second. High PSNR and SSIM scores denote nearly perfect recovery of the original images.
@@
More specifically, with the help of the “zero gradient” module, the aggregated model update $\sum_{i=1}^{n}\delta_{i}$ the attacker obtains is identical to the model update of the victim client $\delta_{target}$, even though the SA protocols are in place. We further argue that the attacker can accurately estimate the gradients from the model update $\delta_{target}$ as it equals local iterations of the gradients [12, 38]. We define the gradients of the first two layers of the target victim as $g_{w_{1}},g_{b_{1}}$ and $g_{w_{2}},g_{b_{2}}$ respectively. The attacker can calculate the following equation to reconstruct the input samples, for $l\in\{1,2,\cdots,k\}$:
$(g_{w_{1}}^{(l+1)}-g_{w_{1}}^{(l)})/(g_{b_{1}}^{(l+1)}-g_{b_{1}}^{(l)}),$ (2)
@@
IV-C Proof of Correctness:
Considering the input $x_{p}$ falls in the $p^{th}$ largest bin, i.e. the brightness of $x_{p}$ denoted by $h(x_{p})$ satisfies $h_{p}<h(x_{p})<h_{p+1}$. Then the following equation holds:
$\frac{g_{w_{1}}^{(p+1)}-g_{w_{1}}^{(p)}}{g_{b_{1}}^{(p+1)}-g_{b_{1}}^{(p)}}$ $=\frac{\nabla_{w_{1}(p+1)}L-\nabla_{w_{1}(p)}L}{\nabla_{b_{1}(p+1)}L-\nabla_{b_{1}(p)}L}$ (3)
@@ IV-C Proof of Correctness:
where $L$ is the loss function, $y$ is the output of the first linear layer, and $\frac{\partial L}{\partial y_{p+1}}=\frac{\partial L}{\partial y_{p}}$ because we let the row vectors of the $w_{2}$ matrix identical. This equation implies $x_{p}$ is perfectly recovered from the gradients of the first linear layer. Because $\mathbf{H}$ covers the whole distribution range of the brightness feature, each input image will fall into one bin and thus can be recovered in this way as long as the image number is smaller than $k$.
V Evaluation
TABLE I: The reconstruction performance of MediLeak over different datasets and reconstruction batch sizes. The rate (sample recovery rate) is on a scale of 1.00
@@ V Evaluation
COVIDx CXR-4 224x224 0.95 120.795 0.99 6.022
200 ChestMNIST(pneumonia) 28x28 0.96 102.722 0.99 0.936
COVIDx CXR-4 224x224 0.89 114.98 0.99 7.003
@@
COVIDx CXR-4 224x224 0.81 95.86 0.99 9.763
@@
Figure 4: The attack performance comparison between MediLeak with other model inversion attacks.
V-A Experimental Settings:
@@
We implemented MediLeak on the PyTorch platform. We ran all the experiments on a server equipped with an Intel Core i7-12700K CPU 3.60GHzX12, one NVIDIA GeForce RTX 3080 Ti GPU, and Ubuntu 20.04.6 LTS.
We use three evaluation metrics including the recovery rate, the peak signal-to-noise ratio (PSNR) score, and the structural similarity index measure (SSIM) score, following the convention of the existing works [20, 21, 22, 23, 24] to evaluate our attack. More specifically, the successful recovery of samples is measured by observing the PSNR and SSIM scores between the original input samples and the reconstructed ones, and checking whether those scores exceed a certain threshold $th$. In our work, we choose $th=20$ for PSNR and $th=0.9$ for SSIM because these thresholds are enough to ensure that the recovered images are visually clear for the attacker to extract all meaningful content from them.
We consider the FL system to have 5 clients with one of them being the attack target per training round and each client conducts 5 local iterations before generating individual model updates. We randomly select $10\%$ of the training set as the auxiliary dataset and aim to recover samples in the test set, which have no intersection with the auxiliary dataset. We assume the test set is partitioned and owned by the 5 clients locally to serve as the local datasets.
@@
We choose the ChestMNIST dataset from the MedMNIST package [29] and the COVIDx CXR-4 dataset [30] as our experiment datasets. The ChestMNIST dataset comprises frontal view X-ray images (1×28×28) of 30805 unique patients with 14 disease labels and we selected data samples related to pneumonia to conduct our experiments, including 78468 training samples and 22433 testing samples. The COVIDx CXR-4 dataset also consists of frontal view chest X-ray images with higher dimensions (resized to 1×224×224) and labels about whether the patient is COVID-positive. The training set contains 67863 samples and the testing set contains 8482 samples.
(a) Recovery Rates.
(b) PSNR Scores.
(c) SSIM Scores.
Figure 5: The performance of MediLeak over different $k$ numbers. $k$ refers to the neuron number of the first linear layer.
@@
TABLE II: Downstream binary classification task on the COVID dataset with a pre-trained (with CXR-3 dataset) ViT model. TPR: True Positive Rate, TNR: True Negative Rate, ACC: Accuracy, AUC: Area Under Receiver Operating Characteristic curve, AUPR: Area Under Precision Recall curve.

### v2

@@ 3. Threat Model
In Fig.1 we demonstrate the threat model of our attack. We consider the parameter server $S$ to be a malicious party that is curious about the training samples held by clients (e.g., private patient images or text records). We consider the parameter server to be a proactive attacker and can actively modify model parameters and architectures from $G$ to $\hat{G}$ to achieve the attack goals, following the same assumption as (Pasquini et al., 2022; Fowl et al., 2021; Zhao et al., 2023; Shi et al., 2025). We assume the communication channels between the server and clients are secured and all messages can be authenticated. We assume the state-of-the-art SA protocols are in place and the server can only get access to the aggregated model updates $\sum_{i=1}^{n}\delta_{i}$ without knowing anything about the individual values $\delta_{i}$. We consider the attacker is able to collect or obtain an auxiliary dataset $D_{aux}$ that has the same data format and can represent the target dataset well. The attacker can leverage various online resources such as publicly available datasets, image searching tools, and image generative tools to fulfill this requirement. For our use case, the availability of public chest X-rays and medical text datasets makes this task trivial. The goal of the attacker is to recover the local data samples of a target client $c_{target}$ just from the aggregated model updates $\sum_{i=1}^{n}\delta_{i}$. This can be mathematically expressed as: $D_{target}=Reverse(\sum_{i=1}^{n}\delta_{i},\hat{G})$.
4. Attack Method
4.1. Attack Overview
As we assume the attacker only possesses the already-masked aggregated model updates $\sum_{i=1}^{n}\delta_{i}$ under the protection of the SA protocol, it is very challenging to identify an end-to-end method that directly reverses the aggregated results back to local training samples $D_{target}$. To address this, we decompose the complex recovery problem into two different distinct tasks including the individual model update retrieval and efficient model update reversion tasks. The first task aims to retrieve the individual model update of the victim from the aggregated model updates, thus breaking the SA protocol. The second task aims to reverse the individual model update back to local samples, achieving the recovery sub-problem.
@@ 4.1. Attack Overview
To accomplish them, we require the attacker to place an additional two-linear-layer module $L_{adv}$ with the rectified linear unit (ReLU) activation function in between at the beginning of the original global model, i.e. $\hat{G}=G\oplus L_{adv}$. The dimension of this module is identical to the image dimension. We require the attacker to initialize the parameters of the linear module $L_{adv}$ to form different adversarial modules and distribute them to different clients accordingly. For all the other clients except the victim, the attacker crafts the “zero gradient” modules $L_{zero}$ to ensure that the gradients and model updates of these clients are always zero, i.e. $\hat{G}_{others}=G\oplus L_{zero}$. By doing this, the attacker guarantees that only the model update of the victim client is exposed, accomplishing task one. For the victim client, the attacker crafts a “linear leakage” module $L_{linear}$, aiming to reverse its model update back to local training samples efficiently, i.e. $\hat{G}_{target}=G\oplus L_{linear}$. This module requires an auxiliary dataset to help generate essential parameters and can ensure that samples are perfectly recovered with a mathematical proof, accomplishing task two. Detailed attack flow and module designs are introduced in the next section.
4.2. Detailed Attack Flow
@@
We demonstrate the attack flow in Fig.2. MedLeak is a two-phase attack including 1) attack preparation and 2) sample reconstruction phases. In the attack preparation phase, the attacker crafts the adversarial global model $\hat{G}$ and initializes it with different model parameters including $L_{zero}$ and $L_{linear}$ (step ①) before publishing them to different clients (step ②). Then in the second phase, the attacker collects the aggregated model updates and uses an analytical method to reverse it back to local training samples (step ③). MedLeak can be used to reconstruct both medical image and non-image data. In the following parts, we will first introduce the overall attack flow and detailed design components via the image data reconstruction task, and then introduce how MedLeak can be customized to accomplish the text reconstruction task.
@@
In this phase, the attacker crafts both the “linear leakage” module and “zero gradient” module once at the outset. Both modules require the estimation of essential parameters of a representative auxiliary dataset $D_{aux}$. In detail, the attacker first estimates the cumulative density function (CDF) of the brightness feature $h(x)$ of the auxiliary dataset $D_{aux}$, denoted by $\psi(h(x))$, to represent the CDF of the local training dataset (which is unavailable), where $x$ refers to the input vector. After that, the attacker divides the distribution $\psi$ into equally $k$ bins by calculating $h_{j}=\psi^{-1}(j/k)$ where $j\in\{1,2,\cdots,k\}$, $\psi^{-1}$ refers to the inverse function of $\psi$, and $k$ equals to the neuron number of the first linear layer. By doing so, the brightness of a random input vector $x$ denoted by $h(x)$ will have the same probability of falling into each bin. This bin vector $\mathbf{H}=\left[h_{1},h_{2},\cdots,h_{k}\right]$ is the key vector to form both attack modules.
@@
Zero Gradient Module: The “zero gradient” module is initialized in the same way as the “linear leakage” module except in step (2), in which the attacker has the bias vector $b_{1}$ equal to $\mathbf{H^{\prime}}=-\left[h_{k},h_{k},\cdots,h_{k}\right]$. By doing so, the output of the first linear layer will always be smaller than zero because $h_{k}$ is the largest possible brightness and all possible input $x$’s brightness is smaller than it. Considering we use the ReLU activation function after the first layer, the input to the second layer and the gradients of the first layer shall always be zero because of the ReLU function’s mathematical property. This results in zero model updates for all clients except the target victim. Therefore, the aggregated model updates are identical to the model update of the victim client, i.e. $\sum_{i=1}^{n}\delta_{i}=\mathbf{0}+\mathbf{0}+\cdots+\mathbf{0}+\delta_{target}=\delta_{target}$, exposing the model update of the target victim.
Figure 3. An example of how the medical text classification system works and where to insert the malicious modules.
@@
More specifically, with the help of the “zero gradient” module, the aggregated model update $\sum_{i=1}^{n}\delta_{i}$ the attacker obtains is identical to the model update of the victim client $\delta_{target}$, even though the SA protocols are in place. We further argue that the attacker can accurately estimate the gradients from the model update $\delta_{target}$ as it equals local iterations of the gradients (McMahan et al., 2017; Shi et al., 2025). We define the gradients of the first two layers of the target victim as $g_{w_{1}},g_{b_{1}}$ and $g_{w_{2}},g_{b_{2}}$ respectively. The attacker can calculate the following equation to reconstruct the input samples, for $l\in\{1,2,\cdots,k\}$:
(2) $(g_{w_{1}}^{(l+1)}-g_{w_{1}}^{(l)})/(g_{b_{1}}^{(l+1)}-g_{b_{1}}^{(l)}),$
@@
4.3. Proof of Correctness
Without the loss of generality, we use $x_{p}$ to denote the local input sample. Considering the input $x_{p}$ falls in the $p^{th}$ largest bin, i.e. the brightness of $x_{p}$ denoted by $h(x_{p})$ satisfies $h_{p}<h(x_{p})<h_{p+1}$. We have the following equation holds:
(3) $\frac{g_{w_{1}}^{(p+1)}-g_{w_{1}}^{(p)}}{g_{b_{1}}^{(p+1)}-g_{b_{1}}^{(p)}}$ $=\frac{\nabla_{w_{1}(p+1)}L-\nabla_{w_{1}(p)}L}{\nabla_{b_{1}(p+1)}L-\nabla_{b_{1}(p)}L}$
@@ 4.3. Proof of Correctness
where $L$ is the loss function, $y$ is the output of the first linear layer, and $\frac{\partial L}{\partial y_{p+1}}=\frac{\partial L}{\partial y_{p}}$ because we let the row vectors of the $w_{2}$ matrix identical.
This equation implies $x_{p}$ is perfectly recovered from the gradients of the first linear layer. Because $\mathbf{H}$ covers the whole distribution range of the brightness feature, each input image will fall into one bin and thus can be recovered in this way as long as the image number is smaller than $k$.
Figure 4. Recovered examples from the COVIDx CXR-4 dataset, Kaggle Brain Tumor MRI dataset, and MedAbstract dataset. The original images are on the left and the recovered ones are on the right. The text samples are truncated due to space limitations. Recovery failure samples are marked in red rectangles.
4.4. Text Data Reconstruction
The medical text records are usually natural language words with discrete values within the input space. They have different mathematical properties from the medical images which are continuous pixel values in the input space. As a result, the medical natural language processing (NLP) models usually have different model architectures and workflows compared to their vision counterparts. In Fig. 3, we demonstrate the architecture of a text classification model. More specifically, for an input sentence $x$, the model first converts it to a sequence of tokens $(t_{1},t_{2},\cdots,t_{l})$, where $l$ refers to the maximum length of the sentence. Then the tokens are fed into the embedding model to be converted into the word embedding vectors, represented by $z=[e_{1},e_{2},\cdots,e_{l}]^{T}$, where all word embedding vectors $e_{i}$ have the same pre-defined embedding dimension. After that, the word embedding vectors are fed into the classification model to produce the output $o$.
Within this model architecture, our original attack strategy of placing the attack module in the front is not effective because the words are all discrete values and cannot be recovered in the same way as the continuous pixel values. Instead, we insert our attack module $L_{adv}$ including both the ”linear leakage” module $L_{linear}$ for the target victim and the ”zero gradient” module $L_{zero}$ for the others between the embedding layer and the classification model. In this way, according to our previous analysis, the attacker can successfully recover the embedding vectors $z=[e_{1},e_{2},\cdots,e_{l}]^{T}$ with the help of these attack modules. The next question is can the attacker further reverse the embedding vectors i.e. $e_{i}$ back to the tokens $t_{i}$? Fortunately, the answer is yes. This is because the word embedding layer maps the discrete tokens into continuous embedding vectors in the embedding space similar to a “lookup table”; therefore, the attacker can simply select the word within the whole word vocabulary that minimizes the distance between the actual embedding vector and the calculated one to be the original token. In practice, we use a single Softmax layer to achieve this reverse function.
Table 2. The reconstruction performance of MedLeak over different datasets and reconstruction batch sizes. The rate (sample recovery rate) is on a scale of 1.00
@@ 4.4. Text Data Reconstruction
COVIDx CXR-4 224x224 0.951 120.795 0.99 6.022
Kaggle Brain Tumor MRI 224x224 0.962 107.783 0.99 6.308
200 ChestMNIST(pneumonia) 28x28 0.960 102.722 0.99 0.936
COVIDx CXR-4 224x224 0.891 114.982 0.99 7.003
Kaggle Brain Tumor MRI 224x224 0.870 98.6421 0.99 6.738
@@
COVIDx CXR-4 224x224 0.810 95.864 0.99 9.763
Kaggle Brain Tumor MRI 224x224 0.796 92.954 0.99 8.767
Analysis: Similar to the image recovery task, the performance of the text recovery task is largely affected by the number of recovery bins $k$. We assume during the FL training process, $m$ text data samples are organized in a batch with each one having a maximum length of $l$ words. We clarify that the maximum sentence length $l$ does not affect the recovery performance as long as it is not super large (e.g. $10^{4}$), which we will later justify in the experiment section. The recovery performance is still largely determined by the relationship between $m$ and $k$. When $k\geq m$, each embedding vector is perfectly recovered within one bin, which further leads to excellent text recovery performance. But when $k<m$, there will be recovery collisions within certain bins, and the text recovery performance drops.
5. Evaluation
@@
Figure 5. The attack performance comparison between MedLeak with other model inversion attacks.
5.1. Experimental Settings
@@
We implemented MedLeak on the PyTorch platform. We ran all the experiments on a server equipped with an Intel Core i7-12700K CPU 3.60GHzX12, one NVIDIA GeForce RTX 3080 Ti GPU, and Ubuntu 20.04.6 LTS.
We considered the FL system to have 5 clients with one of them being the attack target per training round. We assumed each client would perform 5 local iterations before generating the individual model updates. We randomly selected $10\%$ of the training set as the auxiliary dataset and aimed to recover samples in the test set, which have no intersection with the auxiliary dataset. We assumed the test set is partitioned and owned by the 5 clients locally to serve as the local datasets. For the defense mechanism, we considered the system to be protected by the SA protocol in (Bonawitz et al., 2017). Therefore, we cannot get access to the individual model updates (which have been cryptographically masked) and we launched our attack solely based on the aggregated model updates. Note that our attack can not only break the SA protocol in (Bonawitz et al., 2017) because MedLeak breaks the abstract aggregation function of the SA protocols regardless of their implementation details. All SA mechanisms realizing the function can be broken, including (Bell et al., 2020; Choi et al., 2020; Guo et al., 2020; Kadhe et al., 2020; Pillutla et al., 2022; Burkhalter et al., 2021).
@@
For the image recovery task, we chose the ChestMNIST dataset from the MedMNIST package (Yang et al., 2023), the COVIDx CXR-4 dataset (Wang et al., 2020), and the Kaggle Brian Tumor MRI dataset (Nickparvar, 2022) as our experiment datasets. The ChestMNIST dataset comprises frontal view X-ray images (1×28×28) of 30805 unique patients with 14 disease labels and we selected data samples related to pneumonia to conduct our experiments, including 78468 training samples and 22433 testing samples. The COVIDx CXR-4 dataset also consists of frontal-view chest X-ray images with higher dimensions (resized to 1×224×224) and labels about whether the patient is COVID-positive. The training set contains 67863 images and the testing set contains 8482 images. The Kaggle Brain Tumor MRI Dataset contains 7023 images of human brain MRI images which are classified into 4 classes: glioma - meningioma - no tumor and pituitary. Its training set contains 5712 images and the testing set contains 1311 images.
We used four evaluation metrics including the recovery rate, the attack time, the peak signal-to-noise ratio (PSNR) score, and the structural similarity index measure (SSIM) score, following the convention of the existing works (Zhu et al., 2019; Zhao et al., 2020; Geiping et al., 2020; Yin et al., 2021; Fowl et al., 2021) to evaluate our attack over the image recovery task. More specifically, the successful recovery of samples was measured by observing the PSNR and SSIM scores between the original input samples and the reconstructed ones and checking whether those scores exceed a certain threshold $th$. In our work, we chose $th=20$ for PSNR and $th=0.9$ for SSIM because these thresholds are enough to ensure that the recovered images are visually clear for the attacker to extract all meaningful content from them.
For the text recovery task, we chose the MedAbstract dataset (Schopf et al., 2023) as our experiment dataset. The MedAbstract data set consists of 14438 medical abstracts describing the patients’ health conditions with each one consisting of a few hundred words. The patients’ conditions are classified into 5 different classes including digestive system diseases, cardiovascular diseases, neoplasms, nervous system diseases, and general pathological conditions.
We used three evaluation metrics including the recovery rate, the word error rate (WER), and the attack time to evaluate MedLeak’s performance over the text recovery task. Similar to the vision task, the recovery rate was defined as the ratio between the number of successfully recovered text samples achieving WERs lower than threshold 0.05 and the total sample number. Meanwhile, the WER was calculated as the portion of recovery failure words in text samples. In our experiment, we focused on the WER of the successfully recovered samples to evaluate their recovery quality.
@@
Table 3. Downstream binary classification task on the COVID dataset with a pre-trained (with CXR-3 dataset) ViT model. TPR: True Positive Rate, TNR: True Negative Rate, ACC: Accuracy, AUC: Area Under Receiver Operating Characteristic curve, AUPR: Area Under Precision Recall curve.